OpenStudy (anonymous):

1/2x + 1/4x = 16 a.16 b.21.3 c.32 d.64

OpenStudy (anonymous):

put all the possible answers instead of x and see which is correct?

OpenStudy (anonymous):

wat do you mean

OpenStudy (anonymous):

look the answer is a or b or c or d isnt? it one of them is correct. we know this. so one of them will be correct when you put instead of x

OpenStudy (anonymous):

well i think a is wrong i guess and b?

OpenStudy (anonymous):

dont need to learn any algebra put all the possible answers instead of x

OpenStudy (anonymous):

um can you explain wat r u tlking about i seriously dont understand i am super bad at math

OpenStudy (anonymous):

no i m giving the shortest way you dont need any knowladge to solve. i ll tell basically the question is this x+1=2 what is x and the optinions are these: a1 b 2 c 3 d4 now dont solve question do only this put the value of a b c d instead of x i ll put d and c for you rest 2 optinion you will put and see which will be right for equation now in d if the answer is d the value of x is 4 so 4+1=2 is it true? no so the answer is not d lets think answer is c if the answer is c the value of x is 3 so 3+1=2 is it true? no it is not so the answer is not c. now put b and a instead of x and see which is correct. if you udnerstand this you will be able to solve any question in 3 second withouth any knowladge

OpenStudy (anonymous):

uhhh seriously dont no what are you talking bout but i got in mind dat c and is not the answer right?

OpenStudy (anonymous):

well i think is b so can you help with this question please 6 – x – x = 18

OpenStudy (anonymous):

a.-6 b. 6 c.9 d.-9

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

wat???